A rocket is shot straight up from the earth, with a net acceleration (= acceleration by the rocket engine minus gravitational pullback) of during the initial stage of flight until the engine cut out at sec. How high will it go, air resistance neglected?

1
Expert's answer
2021-11-03T14:27:22-0400

The acceleration of a rocket is given by


"\dfrac{d^2s}{dt^2}=7t"

The initial velocity of rocket at the time "t=0"  is zero


"v(0)=0"

Consider


"v(t)=\dfrac{ds}{dt}"

Then


"\dfrac{dv}{dt}=7t"

"v=\int7tdt"

"v(t)=\dfrac{7}{2}t^2+v(0)"

"v(t)=\dfrac{7}{2}t^2"

"s(t)=\int\dfrac{7}{2}t^2dt"

"s(t)=\dfrac{7}{6}t^3+s(0)"

At "t = 0," the rocket is on the surface of the earth, then


"s(t)=\dfrac{7}{6}t^3+R_E"

The engine cuts out at "t = 10 s" . The velocity at engine cutoff is

"v_1=v(10)=\dfrac{700}{2}\ m/s=350\ m/s"

The distance of the rocket from the center of the earth at engine cutoff is


"s_1=s(10)=\dfrac{7000}{6}\ m+R_E"

"\approx6379267\ m"

In the unpowered phase of flight, only gravity acts on the rocket. So we need to figure out how far up an object that starts at "s=s_1" with velocity "v = v_1" will go.

We have


"v\dfrac{dv}{ds}=-\dfrac{gR_E^2}{s^2}"

"vdv=-\dfrac{gR_E^2}{s^2}ds"

Integrate


"\dfrac{v^2}{2}=\dfrac{gR_E^2}{s}+\dfrac{C}{2}"

"v^2=\dfrac{2gR_E^2}{s}+C"

"C=v_1^2-\dfrac{2gR_E^2}{s_1}"

The rocket will be at it’s highest point (greatest value of s) when it comes to a stop, just before starting to fall back to earth


"v=0=>\dfrac{2gR_E^2}{s}+v_1^2-\dfrac{2gR_E^2}{s_1}=0"

"s=\dfrac{2gR_E^2s_1}{2gR_E^2-v_1^2s_1}"

"s\approx\dfrac{2(9.81\ m/s^2)(6378100\ m)^2(6379267\ m)}{2(9.81\ m/s^2)(6378100\ m)^2-(350\ m/s)^2(6379267\ m)}"

"\approx6385519\ m"

Thus, the rocket reachs about


"6385519\ m-6378100\ m=7419\ m"

above the surface of the earth.